### v1

# PAOLI: Pose-free Articulated Object Learning from Sparse-view Images

Abstract
We present a novel self-supervised framework for learning articulated object representations from sparse-view, unposed images. Unlike prior methods that require dense multi-view observations and ground-truth camera poses, our approach operates with as few as four views per articulation and no camera supervision. To address the inherent challenges, we first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses. A progressive disentanglement strategy further separates static from moving parts, enabling robust separation of camera and object motion. Finally, we jointly optimize geometry, appearance, and kinematics with a self-supervised loss that enforces cross-view and cross-pose consistency. Experiments on the standard benchmark and real-world examples demonstrate that our method produces accurate and detailed articulated object representations under significantly weaker input assumptions than existing approaches.
1 Introduction
Interacting with articulated objects, objects that can change shape or position through joints, is a natural part of everyday human activity. Think of opening a cupboard door, sliding out a drawer, or lifting a lid. These objects are characterized by their ability to move in specific ways, often defined by a set of joints that allow for rotation or translation. Understanding how these objects operate is crucial for tasks such as manipulation, and control in robotics.
Earlier research [35, 13, 39] tackled this challenge by using detailed ground-truth information such as 3D object shapes, articulation parameters, or part segmentations. While effective, this approach depends on high-quality 3D scans and labor-intensive manual annotations, which are both costly and difficult to scale. To address this, recent methods have shifted toward learning directly from raw, unlabeled data. Instead of relying on manual labels, these approaches capture the motion of articulated objects and use it to infer their structure and kinematics [22, 7, 10]. For instance, PARIS [22] and ArticulateYourNeRF [7] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [30] to model 3D geometry and appearance while separating moving parts from static ones. ArticulatedGS [10] improves over [22, 7] by modeling articulated objects through Gaussian Splatting (GS) [15], offering improved efficiency and accuracy. Despite these advances, existing methods still rely on dense observations, requiring around 100 paired views of the same object across two different articulated poses. Such requirements are rarely feasible in practical applications. Moreover, these approaches typically assume known camera poses, an assumption does not hold in many real-world scenarios.
Figure 1: Given few unposed views of an articulated object over different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while also estimating its part segmentation and joint axis of its moving parts (shown as arrows) in 3D. The reconstructed object can be rendered from novel viewpoints and articulated poses.
In this work, we propose a novel approach for learning articulated object representations from sparse-view, unposed images. This setting introduces two major challenges. First, with only a handful of views (e.g., four per articulation) and no camera information, reconstructing the object’s 3D structure becomes highly under-constrained. Second, recovering articulated motion and camera motion from images are tightly coupled, making it difficult to disentangle object motion from viewpoint changes.
To tackle these challenges, we introduce a novel framework, PAOLI, that learns to reconstruct articulated objects from sparse-view images without requiring camera poses. For the first challenge, we build on recent advances in sparse-view 3D reconstruction without camera supervision for static scenes. Specifically, we reconstruct the 3D shape of the object independently for each articulation [53], and then learn a deformation field that establishes dense correspondences across the reconstructed shapes. For the second challenge, we introduce a progressive disentanglement strategy: the procedure first identifies static components of the object before progressively isolating the moving parts. This step-by-step process enables a robust separation of camera motion from articulated motion. Finally, we jointly optimize geometry, appearance, and kinematics using a self-supervised loss that enforces consistency across different views and poses.
We evaluate our method on several benchmark datasets and show that it achieves accurate and detailed articulated object representations from sparse, unposed inputs, and significantly outperforming prior approaches that require dense multi-view data and ground-truth camera poses.
2 Related Work
Pose-free 3D reconstruction
Classical Structure from Motion (SfM) [1, 26, 37, 38] jointly estimates camera poses and scene geometry via image registration, triangulation, and bundle adjustment. Recent learning-based multi-view stereo (MVS) methods [51, 46, 20, 44] achieve high-quality reconstructions without explicit feature matching. DUST3R [46] uses transformers to predict dense point maps and poses from image pairs, though bundle adjustment is still required. MASt3R [20] improves scalability with coarse-to-fine matching for high-resolution inputs, while VGGT [44] extends this paradigm with multi-task learning to jointly predict point maps, poses, depth, and tracks. These methods, however, focus on static scenes and do not explicitly model articulated motion or object parts.
Sparse View Gaussian Splats
Neural rendering [30, 15, 12, 45, 33] enables photorealistic novel view synthesis by rendering from learned 3D representations. More recently, 3D GS [15] has emerged as particularly compelling, offering efficient, high-quality representations [8, 5, 29, 18, 49]. Standard GS requires dense multi-view images with accurate poses. To relax this, some works use depth priors [21, 54, 6] or diffusion guidance [24, 16, 28], while others directly predict Gaussian parameters with feed-forward networks [52, 53, 40, 41, 2, 4]. Splatter Image [40] and LGM [41] handle single images, while MVSPlat [4] and PixelSplat [2] target sparse multi-view with known poses. FreeSplatter [53] and NoPoSplat [52] remove pose supervision by jointly predicting poses and Gaussians, but remain limited to static scenes without articulated motion.
Articulated Object Reconstruction
Reconstructing articulated objects is challenging due to coupled geometry and kinematics. Some methods focus on structure or motion alone [11, 17, 3, 9, 25, 14], while others use generative part-level graphs [23, 19]. More recent works jointly model geometry and appearance through optimization [32, 42, 47, 22, 7, 48, 27, 10]. PARIS [22] reconstructs parts and motions from dense multi-view scans with known poses; ArticulateYourNeRF [7] improves robustness via iterative segmentation and motion learning; while ArticulatedGS [10], ArtGS [27], and DTA [48] enhance quality with GS and depth priors. However, all rely on dense multi-view inputs and calibrated poses. Our approach instead reconstructs articulated objects from sparse, unposed views by leveraging recent feed-forward MVS advances.
3 Method
Assume that we are given two sets of images, each capturing an articulated object from sparse and unknown viewpoints. Within each set, the object remains in a fixed pose, while its articulated pose differs across the sets. Our objective is to reconstruct a digital twin of the object that includes both high-fidelity 3D geometry and appearance, and articulation parameters, enabling physically plausible dynamic rendering.
Our method consists of three main stages. First, we reconstruct the 3D geometry and estimate camera parameters for each set of sparse-view images independently using Gaussian representations. Next, we learn a deformation field to align the geometry between two sets. Based on the estimated deformation, we establish the initial dense correspondences and part segmentation. Finally, we jointly optimize the geometry, part segmentation, and articulation parameters to improve the overall reconstruction quality. An overview of our method is shown in Fig.˜2. Since the optimization of the geometry, part segmentation, and articulation parameters is intertwined and requires good initialization, the first two stages are crucial for the success of the final joint optimization.
Figure 2: Overview of our method pipeline. Given two sets of K-view images (K=4) of an articulated object in different states, our approach reconstructs high-fidelity 3D geometry and estimates articulation parameters through a three-stage process: (1) initialization of Gaussian splats and camera parameters (top left), (2) Gaussian alignment via deformation fields (top right), and (3) joint optimization of geometry and articulation parameters (bottom).
3.1 Initialization of 3D Gaussian Splats
Given two sets of images, a source set $\mathcal{I}^{s}=\{I^{s}_{k}\}_{k=1}^{K}$ and a target set $\mathcal{I}^{t}=\{I^{t}_{k}\}_{k=1}^{K}$, each containing $K$ (e.g.4) pose-free views of the object, we first construct initial 3D Gaussian splats $G^{s}$ and $G^{t}$ independently. Each set $G=\{\mathbf{g}_{i}\}_{i=1}^{N}$ consists of $N$ Gaussian points, where each point $\mathbf{g}=(\bm{\mu},\mathbf{q},\mathbf{s},o,\mathbf{c})$ is parameterized by its mean $\bm{\mu}$, rotation $\mathbf{q}$, scale $\mathbf{s}$, opacity $o$, and color $\mathbf{c}$.
Since the standard 3D GS [15] typically requires many input views for accurate reconstruction, we use FreeSplatter [53], a feed-forward neural network trained to reconstruct 3D Gaussian splats from sparse views. As FreeSplatter is designed for static scenes, we apply it independently to each image set to obtain initial Gaussian splats and camera parameters.
Despite its ability to work in the sparse setting, the Gaussians estimated by FreeSplatter are not accurately pixel-aligned with the input images, resulting in suboptimal camera extrinsic estimates. Additionally, the estimated camera intrinsics often differ between the two sets. Hence, we use their average as the final intrinsic parameters. To improve the pixel alignment, we refine the extrinsic parameters $\Pi^{s}$ and $\Pi^{t}$ by directly optimizing them on the reconstructed Gaussian splats over their respective multi-view set, minimizing photometric loss via rendering. This two-step camera refinement yields better alignment between the estimated camera parameters and the 3D geometry, providing a better initialization for the subsequent steps.
3.2 Gaussian Alignment via Deformation Fields
A central challenge in building digital twins of articulated objects from pose-free images lies in aligning the canonical pose across different observation sets. We frame this alignment as estimating the deformation between the Gaussian representations $G^{s}$ and $G^{t}$. A straightforward strategy is to optimize the mean and rotation of each Gaussian point independently. However, this naive approach breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
To address this, we introduce a feed-forward deformation network, $\mathcal{F}_{\text{deform}}$, which predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$. By constraining the network to generate low-frequency deformations, we enforce local rigidity while still capturing global motion. Specifically, $\mathcal{F}_{\text{deform}}$ maps the positional encoding of Gaussian means to a rigid transformation consisting of a rotation, translation, and scaling factor:
$\{\mathbf{R},\mathbf{t},s\}=\mathcal{F}_{\text{deform}}(\text{PE}(\bm{\mu}))$ (1)
where $\mathbf{R}\in\text{SO}(3)$ is parameterized by Euler angles $\in\mathbb{R}^{3}$ and converted to rotation matrices, $\mathbf{t}\in\mathbb{R}^{3}$ represents the translation vector, $s\in\mathbb{R}$ is a global scaling factor shared across all points, and $\text{PE}(\bm{\mu})$ is the positional encoding of the Gaussian means. This transformation is then applied to the source Gaussian points to obtain the deformed target representation.
The means of the source Gaussian points $\bm{\mu}\in\mathbb{R}^{3}$ are encoded as:
$\text{PE}(\bm{\mu})=[\sin(2^{k_{0}}\pi\bm{\mu}),\cos(2^{k_{0}}\pi\bm{\mu})]$ (2)
where $k_{0}$ represents the base frequency parameter. The parameter $k_{0}$ directly influences the rigidity of the deformation field: smaller values result in lower-frequency encodings that naturally constrain the deformation field to be more rigid, effectively guiding the network’s capacity to learn low-frequency non-rigid transformations.
Deformation Network Training
To optimize the parameters of $\mathcal{F}_{\text{deform}}$, we minimize the Chamfer distance $\mathcal{L}_{\text{CD}}$ between the means of the deformed source Gaussian $\hat{G}^{t}$ and the target one ${G}^{t}$, as well as $L1$ photometric loss $\mathcal{L}_{\text{photo}}$ between their rendered images $\mathcal{R}(\hat{G}^{t}),\mathcal{R}(G^{t})$ respectively:
$\min_{\mathcal{F}_{\text{deform}}}\mathcal{L}_{\text{CD}}(\hat{G}^{t},G^{t})+\mathcal{L}_{\text{photo}}(\mathcal{R}(\hat{G}^{t}),\mathcal{R}(G^{t}))$ (3)
where $\hat{G}^{t}=\{\hat{\mathbf{g}}_{i}^{t}\}_{i=1}^{N^{s}}$ represents the deformed source Gaussians, with each point $\hat{\mathbf{g}}^{t}$ obtained by applying the transformation predicted by $\pi=\mathcal{F}_{\text{deform}}(\bm{\mu})$ to the corresponding source point $\mathbf{g}^{s}\in G^{s}$, where $\pi$ is a similarity transformation containing rotation $\mathbf{R}$, translation $\mathbf{t}$, and scaling $s$. Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq.˜1, and apply the transformation to the means and rotation of the source Gaussian points. With $\hat{\bm{\mu}}^{t}=s\mathbf{R}\bm{\mu}^{s}+\mathbf{t}$ and $\hat{\mathbf{q}}^{t}=\mathbf{q}^{s}\otimes\mathbf{q}(\mathbf{R})$, we can obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, where $\otimes$ denotes quaternion multiplication and $\mathbf{q}(\mathbf{R})$ converts the rotation matrix $\mathbf{R}$ to its quaternion representation. For simplicity, we denote this process as $\hat{\mathbf{g}}=\pi\circ\mathbf{g}$.
3.3 Joint Optimization
To obtain an initial segmentation of object parts and their corresponding rigid transformations, we leverage TEASER [50], a robust pose estimation algorithm for point correspondences. We apply TEASER to the set of correspondences between the deformed source and the target Gaussians. Since TEASER was originally designed for static scenes, it identifies the largest rigidly moving component, typically the part with greatest volume, as the inlier set, while treating the remaining points as outliers.
We then organize these sets into a tree structure: the inlier points form the root node, and the outliers are assigned to leaf nodes. TEASER is applied recursively to the leaf nodes to estimate part-level similarity transformations. This iterative process progressively adds parts to the tree, with the root representing the dominant component and leaves corresponding to smaller articulated parts. For simplicity, we describe the two-part case, though the method naturally extends to multiple parts. The result is a set of similarity transformations, $\pi_{r}$ and $\pi_{l}$, for the root and leaf parts, along with segmentation labels $p^{s}\in\{r,l\}$ for each Gaussian point in the source state.
Next we proceed to the joint optimization of the Gaussian splats, part segmentation, and articulation parameters. We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$ with $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ and treat $\pi_{r}$ and $\pi_{l}$ as learnable parameters to refine the rigid transformation. To obtain the deformed Gaussian splats, we apply the transformations to each Gaussian point based on its segmentation label: $\pi_{r}\circ\mathbf{g}$ if $p^{s}=r$, or $\pi_{l}\circ\mathbf{g}$ if $p^{s}=l$. Since the initial estimation is not perfect, we also apply a displacement term $\delta\bm{\mu}$ to the means of the deformed Gaussian points to compensate for non-rigidity. To simplify the notation, we use an auxiliary function $f(\mathbf{g},\pi,\delta\bm{\mu})$ that first applies the transformation $\pi$ to $\mathbf{g}$ and then applies the displacement $\delta\bm{\mu}$ to the mean of the transformed Gaussian. This yields the deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}|\bar{\mathbf{g}}^{t}=f(\bar{\mathbf{g}}^{s},\pi_{p^{s}},\delta\bm{\mu}),\bar{\mathbf{g}}^{s}\in\bar{G}^{s}\}$, with the $\delta\bm{\mu}$ denoting the displacement for the means.
Since the initial segmentation and and transformation for each part is imperfect. We often end up with detached parts after the deformation. To deal with this, we refine the estimated $\pi_{r}$ and $\pi_{l}$ with the initial segmentation labels in the first phase of optimization using both the Chamfer distance between the transformed $\bar{G}^{t}$ and $G^{t}$, and the rendering loss computed in both the source and target states. We define the joint loss as:
$\mathcal{L}_{\text{joint}}=$ $\mathcal{L}_{\text{CD}}(\bar{G}^{t},G^{t})+$ (4)
$\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),\mathcal{R}(G^{t}))+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),\mathcal{R}(G^{s}))$
and our optimization objective is to minimize the joint loss:
$(\pi_{r}^{*},\pi_{l}^{*})=\arg\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (5)
Note that in this phase, $\bm{\delta}\mu$ is set to zero.
In the second phase of the optimization, we optimize the Gaussian splats to improve the geometric accuracy, especially to handle occlusion between source and target states. We also allow the displacement term $\delta\bm{\mu}$ to be optimized to compensate for non-rigidity caused by imperfect transformation estimation or geometry mismatch, but still penalize the displacement of the Gaussian splats with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$. For better optimization stability suggested in ArticulateYourNeRF [7], we update the segmentation labels heuristically. Newly split or duplicated points inherit the segmentation labels from their parent points. To correct points with incorrect initial segmentation, we check those points with large $\delta\bm{\mu}$, which indicates high non-rigidity, and update their segmentation labels following this rule:
$p=\arg\min_{p\in\{r,l\}}\|\bar{\bm{\mu}}^{p}-\bar{\bm{\mu}}^{t}\|_{2}$ (6)
with $\bar{\bm{\mu}}^{p}$ for the means of $\bar{\mathbf{g}}^{p}=\pi_{p}\circ\bar{\mathbf{g}}^{s},p\in\{r,l\}$. And in this phase, our optimization objective is to minimize the joint loss with the regularization term:
$\min_{\{\delta\bm{\mu}\},\bar{G}^{s},\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}+\mathcal{L}_{\text{rigid}}$ (7)
and the update of the segmentation labels is:
$\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (8)
In the final phase, we focus on improving visual quality and articulation estimation. We update the $\pi_{r}$, $\pi_{l}$, and $\mathbf{c}$ of the Gaussian splats without the displacement term $\delta\bm{\mu}$ for pure rigid transformation supervised by photo loss with the original input images as:
$\min_{\{\pi_{r},\pi_{l},\mathbf{c}\}}\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),I^{s})+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),I^{t}).$ (9)
In practice, we use 500 iterations for the first phase, 2500 iterations for the second phase, and 2000 iterations for the final phase. After 5000 iterations at total, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
3.4 Extension to Multiple Moving Parts
Figure 3: Illustration of the extension to multiple parts. We progressively estimate the segmentation and articulation one part at a time with extra images for new target articulation states.
For simplicity, we have described our method for two parts, but it can be easily extended to objects with multiple parts as shown in Fig. 3. This extension requires additional input images for the new moving part which are used to initialize the 3D Gaussian splats and camera parameters for each set of images in §3.1.Here we use the Gaussian splats from the second row of input images as source Gaussian and the first row as target Gaussian shown in Fig. 3. We first go through the pipeline with Gaussian deformation in §3.2 and ¢3.3 to model the motion of the first part, which is the light blue arm shown in Fig. 3. After the optimization, we obtain the articulation parameters and segmentation of the first arm. Based on this, we use the same source Gaussian together with the Gaussian in the bottom row as target Gaussian to model the second part by repeating deformation §3.2 and joint optimization §3.3. Note that there is always only 1 articulated part between source and target Gaussian.
4 Experiments
[…]
Motion Ang Err $\downarrow$ AGS[10] GT 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
AYN[7] GT 52.986 60.834 2.105 42.101 81.157 88.316 74.938 24.850 failed 66.718
AGS[10] VGGT 49.001 66.191 82.359 51.069 21.378 40.371 288.473 83.680 67.020 71.801
Ours 0.095 0.914 1.020 2.365 0.839 1.337 2.905 1.737 0.878 0.525
Pos Err $\downarrow$ AGS[10] GT 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
AYN[7] GT 0.021 0.073 0.033 2.116 0.221 0.231 0.916 0.334 - -
AGS[10] VGGT 0.455 0.254 0.033 0.775 0.741 0.090 0.412 0.513 - -
Ours 0.180 0.007 0.106 0.042 0.058 0.196 1.042 0.157 - -
Motion Dist $\downarrow$ AGS[10] GT 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
AYN[7] GT 122.730 562.642 10.115 20.583 111.252 118.564 50.044 81.205 failed 0.373
AGS[10] VGGT 104.417 54.799 43.799 45.248 27.312 57.388 53.142 34.177 0.202 0.297
Ours 0.861 2.279 0.788 2.654 1.516 0.930 2.227 2.926 0.017 0.118
Rendering PSNR $\uparrow$ AGS[10] GT 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
AYN[7] GT 10.827 14.631 13.411 15.540 14.675 15.826 16.801 15.784 failed 14.018
AGS[10] VGGT 10.740 16.403 17.737 15.123 18.755 14.866 12.990 16.082 18.268 18.483
Ours 24.659 26.920 26.491 23.434 28.823 27.900 27.016 25.408 29.041 25.493
SSIM $\uparrow$ AGS[10] GT 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
AYN[7] GT 0.718 0.893 0.852 0.816 0.904 0.881 0.855 0.889 failed 0.814
AGS[10] VGGT 0.693 0.861 0.883 0.801 0.910 0.890 0.821 0.878 0.905 0.874
Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
Geometry CD $\downarrow$ AGS[10] GT 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
AYN[7] GT 677.761 164.321 124.728 335.108 258.443 125.208 740.898 1575.325 failed 62.890
AGS[10] VGGT 476.935 48.592 77.646 26.251 21.667 112.274 632.828 24.736 26.769 15.599
Ours 0.856 2.051 0.610 6.449 0.264 1.216 8.935 1.615 1.788 6.443
Table 1: Comparisons to the baseline methods with 4-view images from each articulation states (source and target). Objects with $\dagger$ are with prismatic joints.
4.1 Experimental Setup
Dataset.
We evaluate our method on the synthetic 3D PartNet-Mobility dataset [31], following the evaluation protocol established by previous work [22, 7, 10]. Our evaluation metrics encompass three key aspects: (1) For rendering, we use PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index) to evaluate the quality of the rendered images for unseen views in both source and target articulation states. (2) For articulation parameter estimation, we use the angular error, position error of the estimated motion axis and additional motion distance to evaluate the accuracy of articulation parameters as in previous work [7, 22, 10]. More details are provided in the supplementary material. (3) For 3D reconstruction precision, we use the Chamfer Distance (CD) to evaluate the precision of the estimated 3D geometry. For PSNR, SSIM and CD, we report the average values of the two articulation states.
Baselines.
To the best of our knowledge, no existing work addresses self-supervised articulated object modeling from few-shot unposed images. Nevertheless, we can evaluate our approach by comparing it with state-of-the-art methods that utilize posed images for articulated object modeling, specifically ArticulatedGS (AGS) [10]. For fair comparison, we employ VGGT [44] to estimate camera poses for the few-shot input images and point clouds initialization. We thus compare our method against two baselines: AGS with ground truth poses and AGS with VGGT estimated poses plus VGGT estimated point clouds, using 4 images per articulation state.
[…]
Figure 4: Qualitative evaluation for novel view synthesis in target state. We can see from the results that both AGS-GT and AGS-VGGT fail to reconstruct the object in the setting of 4-view images. In the meanwhile, our method demonstrates similar rendering quality compared to AGS-Full, which is trained with 100 images per articulation state with ground truth camera poses.
4.2 Comparison with Baselines
Table 1 presents the quantitative comparison between our method and the baselines. The results clearly show that existing methods cannot handle the few-shot unposed setting—all three methods perform catastrophically across all metrics. Despite using VGGT-estimated point clouds for initialization, AGS-VGGT still fails to produce meaningful reconstructions. Our method, however, consistently outperforms the baselines across all metrics and object types. The baseline methods exhibit very high Chamfer Distance errors, indicating poor geometry reconstruction quality even at the source state. The poor performance stems from several fundamental limitations: limited input views severely hinder baseline reconstruction capabilities. Since these baseline methods heavily rely on precise 3D reconstruction and camera poses to identify articulated parts, the poor geometry reconstruction quality inevitably leads to inaccurate articulation parameter estimation and degraded novel articulation synthesis quality. As for the AYN method, it raises an error when running on the "Blade" object as it failed to initialize part segmentation due to poor reconstruction quality.
Despite outperforming the baselines, certain objects remain challenging for our method. For instance, the part segmentation of the "Blade" object shows limited accuracy, as illustrated in Fig. 5. We observe that only the blade tip and the blue handle are properly segmented, while the main blade body is classified as the root part. This occurs because the blade’s movement is too small, the pipeline could produce similar visual results during rendering by just moving the blade tip and the blue handle. Additionally, our method experiences performance degradation when dealing with thin articulated parts, like the lid of "Washer" and the door of "Oven". Since reconstructing thin objects is inherently challenging, the resulting geometry errors create larger margins for articulation estimation errors. Besides, objects with large occluded parts are difficult to to maintain consistency in the reconstructed geometry between source and target states. One example would be the USB, which has a large occluded part at target state. Such geometry inconsistency results in performance degradation.
We further examine the baseline failures through visual analysis. Figure 4 shows view synthesis results, where baseline methods generate heavily corrupted renderings with substantial noise—the objects become barely recognizable. Surprisingly, our method matches the quality of AGS-Full, which uses 100 images per articulation state with ground truth poses. Figure 5 displays the 3D Gaussian point distributions with part segmentation. The baselines produce scattered floating points and fragmented geometries, while our method creates coherent 3D structures with proper part segmentation and articulation parameter estimation.
[…]
rowsep=0pt,colsep=1pt RGB input AGS-GT AGS-VGGT AGS-Full Ours GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visualization of 3D Gaussian points with part-level segmentation and articulation axes. Ground truth axises are shown in green, predictions in red. Better view in color and zoom in.
[…]
AGS GT 34.914 0.988 1.279 0.101
VGGT 28.860 0.975 1.367 1.623
AYN GT 32.219 0.962 0.946 0.654
Ours - 28.870 0.975 1.399 1.760
Table 2: Comparative analysis for 16 input views per articulation state.
Figure 6: PSNR w.r.t. number of views per articulation.
[…]
rowsep=1pt,colsep=1pt 4-view 8-view 16-view AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AYN-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative evaluation for the effect of the number of input images. We show novel view synthesis for source state (left) and target state (right) at each settings. From top to bottom: AGS-GT, AYN-GT, AGS-VGGT, Ours.
[…]
rowsep=0pt,colsep=1pt $k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$ Refer to caption Refer to caption Refer to caption Coarse segmentation at source state Refer to caption Refer to caption Refer to caption Deformation Overlay of Gaussian Points
Figure 8: Ablation study for the choice of $k_{0}$ for the deformation field. For segmentation, different color denotes different parts identified using TEASER at source state. For deformation overlay, red points denote the deformed Gaussian points $\hat{G}^{t}$, and green points denote the Gaussian points at target state $\hat{G}^{t}$.
[…]
rowsep=0pt,colsep=1pt RGB input Novel articulation synthesis Part-level segmentation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Real-world objects. We show the input images (left), novel articulation synthesis results (middle), and part-level segmentation with articulation axes (right). The first three objects are single-part articulated objects, while the last one is a multi-part articulated object. For the multi-part object, we collect three sets of input images per articulation state. The segmentation results are shown in different colors for each part. The green lines denote the estimated articulation axes.
4.3 Ablation Studies
Effect of the number of input images.
With just 4-view images, baseline methods struggle to reconstruct objects even when in static state. We investigate how the number of input views affects performance by varying the views per articulation state. Fig. 6 shows that PSNR for baseline methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses. Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation shows in Fig. 7. Tab. 2 shows the comparison at 16 images per state, where AGS with VGGT poses gets close to our performance. However, AGS still encounter reconstruction failures even with 32 images per state when VGGT provides less accurate poses. We suspect that AGS is more sensitive to pose estimation errors as it heavily depends on 2D image information. Our approach leverages FreeSplatter’s 3D priors and 3D-guided deformation, making it robust to pose estimation variations when poses stay consistent relative to the reconstructed splats.
Effect of $k_{0}$ for deformation field
We find that our method is generally robust to the choice of $k_{0}$ in the deformation field. To analyze its effect, we conduct experiments with different $k_{0}$ values, with qualitative results shown in Fig. 8. A smaller $k_{0}$ enforces near-rigid transformations across the whole object, leading to under-deformation and failure to reach the target state (e.g., right side of the second row). In contrast, a larger $k_{0}$ relaxes rigidity too much, causing over-deformation. Both extremes result in poor initial part segmentation and degraded performance. Empirically, we find that setting $k_{0}$ between $10^{-7}$ and $10^{-9}$ achieves a good balance, preserving rigidity while allowing sufficient flexibility for deformation.
4.4 Real-world Experiments
To demonstrate the generalization capability of our approach, we conduct experiments on real-world articulated objects. We evaluate our method on four different objects: a bottle, a projector, a drawer, and a lego toy, collecting four input images for each articulation state. Prior to processing, we employ the Segment Anything Model (SAM) [36] to remove background from the input images. For objects with multiple articulated parts, such as the lego toy with two movable arms, we collect three sets of input images per articulation state as explained in Sec. 3.4. The experimental results, presented in Fig. 9, demonstrate that our method successfully generates seamless reconstruction results for real-world articulated objects, validating its practical applicability beyond synthetic datasets.
5 Conclusion
We introduced PAOLI, a self-supervised framework for learning articulated object representations from sparse, unposed images. Unlike prior methods that require dense multi-view data and known camera poses, PAOLI operates with only a few views per articulation and no camera supervision. By combining sparse-view 3D reconstruction, a deformation field for cross-pose correspondences, and a progressive disentanglement strategy, our method jointly optimizes geometry, appearance, and kinematics while robustly separating camera and object motion.
Our work assumes articulated objects with rigid parts, cannot handle non-rigid deformations. Handling such transformations requires more complex deformation models. Second, we assume that the observations from two different poses of the same object are available. It would be valuable to pursue an active vision approach that allows the system to interact with the object while learning its articulated structure, which could improve robustness and generalization (e.g., [34]). Additionally, our current method relies on comprehensive object coverage with four input views per articulation state; insufficient coverage may degrade reconstruction quality and lead to failure. Integrating diffusion models to synthesize additional viewpoints could offer a promising solution for scenarios with limited view coverage. Finally, temporal information could be leveraged to improve the understanding of articulated motion dynamics.
Supplementary Material
6 Supplementary Material
6.1 Detailed Discussion of Related Work
Here we provide more detailed discussion of the most related articulated object learning work including, PARIS [22], AYN [7] and AGS [10]. As explained in the submission, these techniques assume dense views of the object across two articulation state along with the camera information, unlike our method which is pose-free and works with sparse views. Below, we explain the differences in methodology in more detail.
All the methods including ours uses the synthesis by analysis approach that deforms the object shape in the source articulation state to the target articulation state while learning the geometry, appearance and part segmentation as the intermediate representations. PARIS jointly optimizes the geometry, appearance and articulation parameters from the input images of both source and target states with photometric loss. It employs two separate NeRFs to model ‘static’ and ‘dynamic’ parts of the object. Although the joint optimization is straightforward, it is unstable and fails to obtain accurate results when the initialization of these parameters are noisy, as reported in both AYN and AGS. In addition, PARIS is limited to modeling one moving part only.
Instead, both AYN and AGS follows a multi-step approach, hence reconstructs the geometry and appearance of the object first in the source state. Once the initial geometry and appearance are learned, they use a decoupled optimization for part segmentation and motion parameter estimation that provides more stable optimization. AYN first extracts evenly spaced point cloud from NeRF scans, and then projects the point clouds back on image space. By comparing the difference in the alpha channel between the projected images and the target state images, AYN initializes the part segmentation label. With the initialized labels, it optimizes the part motion parameter so that the projected images with transformed part-points match the image of corresponding target images in alpha channel measured by 2D chamfer distance. Later, once the motion parameters are fixed, and AYN learns a segmentation field to refine the segmentation by minimizing the photo metric loss between rendered images and target images.
Similarly, AGS tackles the instability problem with multi-step approach. To first obtain the segmentation labels for gaussians, by employing a deformation field to learn the displacement for each Gaussian and rotation to correct the quaternions for color rendering. That’s $\hat{\bm{\mu}}=\bm{\mu}+\delta\bm{\mu}$, $\hat{\bm{q}}=\bm{q}\otimes\delta\bm{q}$. This steps aims to obtain the deformed 3D Gaussian that could render similar images in the target states with the corresponding camera poses. The deformation is supervised by photometric loss and a rigidity regularization term. Since the camera poses are aligned with the root part of the objects, the gaussian points in the root part will be static with small displacement $\delta\bm{\mu}$. AGS assigns Gaussians with small displacement to the static part, and the ones with large deformations to the moving part. After freezing their segmentation labels, it estimates a rigid transformation for the dynamic part with chamfer distance w.r.t. the deformed Gaussian (by deformation field) and photometric loss w.r.t. the images in the target states.
However, both AYN and AGS rely on known camera poses that are aligned at the root part between source and target states. This requirement prevents them from initializing part segmentation when camera poses are unknown across different articulation states.
Unlike previous methods that identify parts based on rendering differences (which only relies on 2D information from images) and assume a static root part, our method establishes correspondence across two articulation states directly in 3D space by deforming the 3D Gaussians from the source to the target state. Since $G^{s}$ and $\hat{G}^{t}$ represent the same set of Gaussians with $\hat{G}^{t}$ being the transformed version of $G^{s}$, we can naturally establish matching pairs between them. The deformation is supervised by 3D chamfer distance and photometric loss, ensuring that $\hat{G}^{t}$ closely approximates $G^{t}$. This makes the matching pairs from $G^{s}$ to $\hat{G}^{t}$ effectively equivalent to matching pairs between $G^{s}$ and $G^{t}$. With these correspondences established, we can separate parts and estimate per-part transformations using TEASER for initialization. To further improve rigidity during deformation and enhance performance, we employ positional encoding on the Gaussian means before feeding them to the deformation network.
With the initialization, we can now continue for fine-level optimization. For stability purpose, we also decouple the updates for part segmentation from the joint optimization. We split the optimization into 3 phases. The first phase we try to refine the coarse motion estimation from TEASER. In the second phase, we update part segmentation based on heuristic rules explained in §3.3 in the main paper to refine the segmentation. At the last phase, we optimize the appearance to improve the rendering quality.
6.2 Evaluation Metrics
As we mentioned in the main paper, we will decompose the relative motion between leaf part and root part into axis $\bm{a}$, pivot $\bm{p}$ and angle $\alpha$ for revolute joints and axis $\bm{a}$ and magnitude $d$ for prismatic joints. Since our method is a pose-free method, we need to recover the reconstructed geometry in the ground truth coordinate system. The coordinate transformation can be computed by aligning the position of the estimated camera pose and the ground truth camera pose using Umeyama’s algorithm [43]. To further mitigate the effects of noisy pose estimation, we optimize the transformation matrix initialized with Umeyama’s algorithm by minimizing the chamfer distance between the reconstructed and ground truth geometries. We apply identical preprocessing procedures when evaluating pose-free baseline methods.
Ang Err:
The angular error computes the orientation difference between the predicted axis direction and the ground truth.
Pos Err:
The position error computes the minimum distance between the rotation axis and the ground truth which takes the position of the pivot point into account. Note that this error only applies to revolute joints.
Motion Dist:
For motion distance computes the difference in movements between estimated motion and ground truth. For revolute joints, we compute the geodesic distance between the predicted rotation matrix and the ground truth rotation matrix. For prismatic joints, we compute the distance between the predicted translation magnitude and the ground truth translation magnitude.
6.3 Additional Visualizations
Source Input Target Input Novel View/Articulation Synthesis
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v2

@@
 Abstract
-We present a novel self-supervised framework for learning articulated object representations from sparse-view, unposed images.
-Unlike prior methods that require dense multi-view observations and ground-truth camera poses, our approach operates with as few as four views per articulation and no camera supervision.
-To address the inherent challenges, we first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses.
+We present a methodology to model articulated objects using a sparse set of images with unknown poses.
+Current methods require dense multi-view observations and ground-truth camera poses.
+Our approach operates with as few as four views per articulation and no camera supervision.
+Our central insight is to first solve a robust correspondence and alignment problem between unaligned reconstructions, before part motions can be analyzed.
+We first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses.
 A progressive disentanglement strategy further separates static from moving parts, enabling robust separation of camera and object motion.
-Finally, we jointly optimize geometry, appearance, and kinematics with a self-supervised loss that enforces cross-view and cross-pose consistency.
+Finally, we optimize geometry, appearance, and kinematics jointly with a self-supervised loss that enforces cross-view and cross-pose consistency.
 Experiments on the standard benchmark and real-world examples demonstrate that our method produces accurate and detailed articulated object representations under significantly weaker input assumptions than existing approaches.
 1 Introduction
-Interacting with articulated objects, objects that can change shape or position through joints, is a natural part of everyday human activity.
-Think of opening a cupboard door, sliding out a drawer, or lifting a lid.
-These objects are characterized by their ability to move in specific ways, often defined by a set of joints that allow for rotation or translation.
-Understanding how these objects operate is crucial for tasks such as manipulation, and control in robotics.
-Earlier research [35, 13, 39] tackled this challenge by using detailed ground-truth information such as 3D object shapes, articulation parameters, or part segmentations.
+Automating human capabilities, such as reasoning about and interacting with pivoting doors, sliding drawers, opening boxes, remains an open problem.
+Practical solutions for modeling such articulated objects—rigid objects coupled via joints that allow rotation or translation— from video is beneficial to downstream applications such as robotic manipulation.
+Current approaches require dense, multi-view images across articulation poses, making them impractical.
+In this work, we show that articulated object modeling is possible with only four unposed images per articulation state, vastly reducing capture requirements.
+Early approaches [42, 18, 48] rely on rich supervision such as 3D object shapes, articulation parameters, or part segmentations.
 While effective, this approach depends on high-quality 3D scans and labor-intensive manual annotations, which are both costly and difficult to scale.
-To address this, recent methods have shifted toward learning directly from raw, unlabeled data.
-Instead of relying on manual labels, these approaches capture the motion of articulated objects and use it to infer their structure and kinematics [22, 7, 10].
-For instance, PARIS [22] and ArticulateYourNeRF [7] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [30] to model 3D geometry and appearance while separating moving parts from static ones.
-ArticulatedGS [10] improves over [22, 7] by modeling articulated objects through Gaussian Splatting (GS) [15], offering improved efficiency and accuracy.
-Despite these advances, existing methods still rely on dense observations, requiring around 100 paired views of the same object across two different articulated poses.
-Such requirements are rarely feasible in practical applications.
-Moreover, these approaches typically assume known camera poses, an assumption does not hold in many real-world scenarios.
-Figure 1: Given few unposed views of an articulated object over different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while also estimating its part segmentation and joint axis of its moving parts (shown as arrows) in 3D.
-The reconstructed object can be rendered from novel viewpoints and articulated poses.
-In this work, we propose a novel approach for learning articulated object representations from sparse-view, unposed images.
-This setting introduces two major challenges.
-First, with only a handful of views (e.g., four per articulation) and no camera information, reconstructing the object’s 3D structure becomes highly under-constrained.
-Second, recovering articulated motion and camera motion from images are tightly coupled, making it difficult to disentangle object motion from viewpoint changes.
-To tackle these challenges, we introduce a novel framework, PAOLI, that learns to reconstruct articulated objects from sparse-view images without requiring camera poses.
-For the first challenge, we build on recent advances in sparse-view 3D reconstruction without camera supervision for static scenes.
-Specifically, we reconstruct the 3D shape of the object independently for each articulation [53], and then learn a deformation field that establishes dense correspondences across the reconstructed shapes.
-For the second challenge, we introduce a progressive disentanglement strategy: the procedure first identifies static components of the object before progressively isolating the moving parts.
-This step-by-step process enables a robust separation of camera motion from articulated motion.
-Finally, we jointly optimize geometry, appearance, and kinematics using a self-supervised loss that enforces consistency across different views and poses.
-We evaluate our method on several benchmark datasets and show that it achieves accurate and detailed articulated object representations from sparse, unposed inputs, and significantly outperforming prior approaches that require dense multi-view data and ground-truth camera poses.
+Recent methods reduce supervision by learning from unlabeled multiple views; rather than relying on manual labels, these approaches infer object structure and kinematics from observed motion [30, 10, 16].
+For instance, PARIS [30] and ArticulateYourNeRF [10] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [38].
+These models jointly recover 3D geometry and appearance while separating moving parts from static ones.
+ArticulatedGS [16] further improves efficiency and accuracy by replacing NeRF with Gaussian Splatting (GS) [21].
+Despite these advances, the state of the art still relies on dense, paired observations, requiring around 100 paired views of the same object across two articulation states.
+Such dense capture is rarely practical.
+Moreover, these methods assume that all capture lie in a shared, pre-calibrated coordinate system, i.e., that the object’s static base is perfectly aligned across articulation states via known camera poses.
+In practice, this level of calibration is costly and rarely achievable, severely limiting deployment.
+Figure 1: Given few unposed views of an articulated object across different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while estimating part segmentation and joint axis of its moving parts (shown as arrows) in 3D.
+The reconstructed object can be rendered from novel viewpoints and articulation configurations.
+We instead study articulated object reconstruction and modeling from unposed and sparse views, significantly more realistic yet fundamentally harder setting.
+Traditional 3D reconstruction methods such as structure-from-motion (SFM) [45, 46] or neural-based methods [53, 55] assume static scenes and cannot directly handle object motion.
+Treating each articulation state independently bypasses this issue but produces reconstructions in different coordinate frames, creating a severe misalignment problem.
+Articulated reasoning becomes impossible without first establishing a reliable 3D-to-3D correspondence between these unaligned shapes.
+A common approach is to match sparse keypoints.
+However, keypoints are unreliable under large articulations, on textureless surfaces, and provide insufficient coverage for dense part segmentation.
+This makes them unsuitable for the experimental setting we target.
+We introduce PAOLI, a methodology that learns dense correspondences via a optimized deformation field.
+Leveraging sparse view 3D reconstruction [62, 14], we first obtain independent GS reconstructions for each articulation state.
+A dense deformation field is optimized to warp the source model into alignment with the target state by enforcing 3D geometric consistency and 2D photometric appearance.
+Dense correspondences between the original source model and the target model are then established by using 1-Nearest-Neighbor (1-NN) to match the deformed source model to the target model.
+Finally to separate rigid parts, we introduce an iterative robust estimation algorithm that clusters inliers from the correspondence set using a pose estimator [59], thereby isolating static and moving components and recovering their kinematic parameters.
+In summary, our main contributions are:
+  1. 1.
+    Problem formulation: We are the first to address articulated object reconstruction from sparse, unposed images, identifying the 3D-to-3D correspondence problem between unaligned reconstructions as the central challenge.
+  2. 2.
+    Dense correspondence via deformation: We propose a robust correspondence detector, using a dense deformation field, which is optimized with 3D and 2D losses.
+This approach outperforms keypoint-based correspondence methods and other adapted baselines.
+  3. 3.
+    A new customized solution: We demonstrate the first complete pipeline to successfully recover full articulated models including geometry, parts, and kinematics from sparse, and unposed data.
+These contributions demonstrate that a robust solution to this correspondence problem is achieved not by sparse keypoints or pre-trained models, but by a dense deformation field optimized at test-time for each object, leveraging a combination of 3D geometry and 2D rendering losses.
 2 Related Work
@@
 FreeSplatter [62] and NoPoSplat [61] remove pose supervision by jointly predicting poses and Gaussians, but remain limited to static scenes without articulated motion.
+UFV-splatter [14] improves robustness to image viewpoints by finetuning FreeSplatter with images from unfavorable views, which enhances reconstruction and camera pose estimation.
 Articulated Object Reconstruction
@@
 More recent works jointly model geometry and appearance through optimization [40, 51, 56, 30, 10, 57, 35, 16].
+DITTO [19] as one of the prioneer work that aims to model two-part articulated objects with point cloud input data and assumes aligned coordinate frame with pre-trained feed-forwrd model.
 PARIS [30] reconstructs parts and motions from dense multi-view scans with known poses; ArticulateYourNeRF [10] improves robustness via iterative segmentation and motion learning; while ArticulatedGS [16], ArtGS [35], and DTA [57] enhance quality with GS and depth priors.
@@
 Our approach instead reconstructs articulated objects from sparse, unposed views by leveraging recent feed-forward MVS advances.
+3D-3D Correspondence and Matching
+Finding 3D-to-3D correspondences is a fundamental problem.
+Approaches based on sparse keypoints, from classical descriptors like FPFH [44] to modern learned detectors [28, 64].
+Recent registration methods for 3D Gaussian Splatting (GS) like GaussReg [3], proposed to establish coarse 3D keypoint matching between GS, which still struggles with accurate registration in 3D data, making them unsuitable for dense articulated motion analysis.
+A more robust alternative is to learn dense correspondences via deformation fields [15, 11, 2, 24, 7].
+These approaches, however, typically rely on feed-forward, pre-trained models to predict deformation and are designed for meshes or point clouds, not 3D GS.
+In contrast, our method introduces a test-time, optimization-based framework to learn a dense deformation field specifically for GS, allowing us to solve the articulated motion problem.
+We then use a robust iterative algorithm to segment these dense correspondences into distinct rigid parts and their kinematics.
 3 Method
-Assume that we are given two sets of images, each capturing an articulated object from sparse and unknown viewpoints.
-Within each set, the object remains in a fixed pose, while its articulated pose differs across the sets.
-Our objective is to reconstruct a digital twin of the object that includes both high-fidelity 3D geometry and appearance, and articulation parameters, enabling physically plausible dynamic rendering.
+We address the task of reconstructing articulated objects, defined as a composition of rigid parts connected by 1-DoF revolute or prismatic joints.
+Given two sets of images, each capturing an articulated object from sparse and unknown viewpoints (with a fixed pose within each set but different poses across sets), our objective is to reconstruct a ”digital twin” of the object.
+This model includes high-fidelity 3D geometry, appearance, and articulation parameters, enabling physically plausible dynamic rendering.
 Our method consists of three main stages.
-First, we reconstruct the 3D geometry and estimate camera parameters for each set of sparse-view images independently using Gaussian representations.
-Next, we learn a deformation field to align the geometry between two sets.
-Based on the estimated deformation, we establish the initial dense correspondences and part segmentation.
-Finally, we jointly optimize the geometry, part segmentation, and articulation parameters to improve the overall reconstruction quality.
-An overview of our method is shown in Fig.˜2.
-Since the optimization of the geometry, part segmentation, and articulation parameters is intertwined and requires good initialization, the first two stages are crucial for the success of the final joint optimization.
+First, we reconstruct the 3D geometry and estimate camera parameters for each image set independently.
+Next, we learn a deformation field to align the geometry and establish initial dense correspondences.
+Finally, we jointly optimize the geometry, part segmentation, and articulation parameters.
+An overview is shown in Fig.
+2.
+Since the final optimization is intertwined and requires good initialization, the first two stages are crucial.
 Figure 2: Overview of our method pipeline.
@@
 Each set $G=\{\mathbf{g}_{i}\}_{i=1}^{N}$ consists of $N$ Gaussian points, where each point $\mathbf{g}=(\bm{\mu},\mathbf{q},\mathbf{s},o,\mathbf{c})$ is parameterized by its mean $\bm{\mu}$, rotation $\mathbf{q}$, scale $\mathbf{s}$, opacity $o$, and color $\mathbf{c}$.
-Since the standard 3D GS [15] typically requires many input views for accurate reconstruction, we use FreeSplatter [53], a feed-forward neural network trained to reconstruct 3D Gaussian splats from sparse views.
-As FreeSplatter is designed for static scenes, we apply it independently to each image set to obtain initial Gaussian splats and camera parameters.
-Despite its ability to work in the sparse setting, the Gaussians estimated by FreeSplatter are not accurately pixel-aligned with the input images, resulting in suboptimal camera extrinsic estimates.
-Additionally, the estimated camera intrinsics often differ between the two sets.
-Hence, we use their average as the final intrinsic parameters.
-To improve the pixel alignment, we refine the extrinsic parameters $\Pi^{s}$ and $\Pi^{t}$ by directly optimizing them on the reconstructed Gaussian splats over their respective multi-view set, minimizing photometric loss via rendering.
-This two-step camera refinement yields better alignment between the estimated camera parameters and the 3D geometry, providing a better initialization for the subsequent steps.
-3.2 Gaussian Alignment via Deformation Fields
+Since FreeSplatter is designed for static scenes, we apply it independently to the source set $\mathcal{I}^{s}$ and the target set $\mathcal{I}^{t}$.
+This process yields our two initial 3D Gaussian Splatting models, $G^{s}$ and $G^{t}$, each in its own arbitrary coordinate system.
+This step also provides initial camera parameters (both extrinsics and intrinsics) for each set.
+We then compute a single, averaged intrinsic matrix from the two sets and use this fixed intrinsic parameter for all subsequent optimization steps.
+However, the initial Gaussians and extrinsic estimates are not perfectly pixel-aligned.
+To refine this, we freeze the GS and optimize only the extrinsic parameters ($\Pi^{s},\Pi^{t}$) for each set by minimizing a photometric rendering loss.
+This refinement provides a better initialization for the subsequent steps.
+3.2 Gaussian Correspondences via Deformation Fields
 A central challenge in building digital twins of articulated objects from pose-free images lies in aligning the canonical pose across different observation sets.
-We frame this alignment as estimating the deformation between the Gaussian representations $G^{s}$ and $G^{t}$.
-A straightforward strategy is to optimize the mean and rotation of each Gaussian point independently.
-However, this naive approach breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
-To address this, we introduce a feed-forward deformation network, $\mathcal{F}_{\text{deform}}$, which predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$.
+As argued in our related work, a common approach for recovering pose is to match sparse keypoints [28, 64].
+However, this is often unreliable for dense analysis and struggles with texture-less surfaces.
+A more robust alternative is to find a dense alignment.
+A naive strategy would be to optimize the mean and rotation of each Gaussian point independently, but this breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
+To address this, we learn a per-scene deformation field, parameterized by a lightweight feed-forward network, $\mathcal{F}_{\text{deform}}$.
+This network is optimized from scratch for each object and predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$.
 By constraining the network to generate low-frequency deformations, we enforce local rigidity while still capturing global motion.
@@
 where $\hat{G}^{t}=\{\hat{\mathbf{g}}_{i}^{t}\}_{i=1}^{N^{s}}$ represents the deformed source Gaussians, with each point $\hat{\mathbf{g}}^{t}$ obtained by applying the transformation predicted by $\pi=\mathcal{F}_{\text{deform}}(\bm{\mu})$ to the corresponding source point $\mathbf{g}^{s}\in G^{s}$, where $\pi$ is a similarity transformation containing rotation $\mathbf{R}$, translation $\mathbf{t}$, and scaling $s$.
-Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq.˜1, and apply the transformation to the means and rotation of the source Gaussian points.
+Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq.
+1, and apply the transformation to the means and rotation of the source Gaussian points.
 With $\hat{\bm{\mu}}^{t}=s\mathbf{R}\bm{\mu}^{s}+\mathbf{t}$ and $\hat{\mathbf{q}}^{t}=\mathbf{q}^{s}\otimes\mathbf{q}(\mathbf{R})$, we can obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, where $\otimes$ denotes quaternion multiplication and $\mathbf{q}(\mathbf{R})$ converts the rotation matrix $\mathbf{R}$ to its quaternion representation.
 For simplicity, we denote this process as $\hat{\mathbf{g}}=\pi\circ\mathbf{g}$.
+Pipeline Corr. Foldchair Fridge Laptop Oven Scissor Stapler USB Washer Blade Storage
+Motion Ang Err $\downarrow$ AGS[16] - 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
+Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
+Ours FM-Dense[2] 0.974 13.267 failed failed 5.209 failed 12.603 35.926 3.716 49.799
+Ours Ours 0.095 0.914 1.020 2.365 0.839 1.337 1.872 1.737 0.878 0.525
+Pos Err $\downarrow$ AGS[16] - 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
+Ours GaussReg[3] failed failed failed failed failed failed failed failed - -
+Ours FM-Dense[2] 0.729 0.658 failed failed 0.394 failed 0.281 0.417 - -
+Ours Ours 0.180 0.007 0.106 0.042 0.058 0.196 0.099 0.157 - -
+Motion Dist $\downarrow$ AGS[16] - 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
+Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
+Ours FM-Dense[2] 13.590 25.261 failed failed 3.652 failed 11.285 17.509 0.089 0.192
+Ours Ours 0.861 2.279 0.788 2.654 1.516 0.930 1.877 2.926 0.017 0.118
+Rendering PSNR $\uparrow$ AGS[16] - 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
+Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
+Ours FM-Dense[2] 17.897 19.256 21.161 19.356 23.276 19.002 22.517 20.632 25.741 20.900
+Ours Ours 24.659 26.920 26.491 23.434 28.823 26.532 27.016 27.416 29.041 25.493
+SSIM $\uparrow$ AGS[16] - 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
+Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
+Ours FM-Dense[2] 0.889 0.916 0.923 0.893 0.953 0.930 0.821 0.933 0.970 0.896
+Ours Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
+Geometry CD $\downarrow$ AGS[16] - 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
+Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
+Ours FM-Dense[2] 4.814 6.399 failed failed 0.501 failed 8.935 25.515 2.789 15.599
+Ours Ours 0.856 2.051 0.610 6.449 0.264 1.216 1.430 1.615 1.788 6.443
+Table 1: Comparisons to the baseline methods with 4-view images from each articulation states (source and target). Objects with $\dagger$ are with prismatic joints.
 3.3 Joint Optimization
-To obtain an initial segmentation of object parts and their corresponding rigid transformations, we leverage TEASER [50], a robust pose estimation algorithm for point correspondences.
-We apply TEASER to the set of correspondences between the deformed source and the target Gaussians.
-Since TEASER was originally designed for static scenes, it identifies the largest rigidly moving component, typically the part with greatest volume, as the inlier set, while treating the remaining points as outliers.
-We then organize these sets into a tree structure: the inlier points form the root node, and the outliers are assigned to leaf nodes.
-TEASER is applied recursively to the leaf nodes to estimate part-level similarity transformations.
-This iterative process progressively adds parts to the tree, with the root representing the dominant component and leaves corresponding to smaller articulated parts.
-For simplicity, we describe the two-part case, though the method naturally extends to multiple parts.
-The result is a set of similarity transformations, $\pi_{r}$ and $\pi_{l}$, for the root and leaf parts, along with segmentation labels $p^{s}\in\{r,l\}$ for each Gaussian point in the source state.
-Next we proceed to the joint optimization of the Gaussian splats, part segmentation, and articulation parameters.
-We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$ with $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ and treat $\pi_{r}$ and $\pi_{l}$ as learnable parameters to refine the rigid transformation.
-To obtain the deformed Gaussian splats, we apply the transformations to each Gaussian point based on its segmentation label: $\pi_{r}\circ\mathbf{g}$ if $p^{s}=r$, or $\pi_{l}\circ\mathbf{g}$ if $p^{s}=l$.
-Since the initial estimation is not perfect, we also apply a displacement term $\delta\bm{\mu}$ to the means of the deformed Gaussian points to compensate for non-rigidity.
-To simplify the notation, we use an auxiliary function $f(\mathbf{g},\pi,\delta\bm{\mu})$ that first applies the transformation $\pi$ to $\mathbf{g}$ and then applies the displacement $\delta\bm{\mu}$ to the mean of the transformed Gaussian.
-This yields the deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}|\bar{\mathbf{g}}^{t}=f(\bar{\mathbf{g}}^{s},\pi_{p^{s}},\delta\bm{\mu}),\bar{\mathbf{g}}^{s}\in\bar{G}^{s}\}$, with the $\delta\bm{\mu}$ denoting the displacement for the means.
-Since the initial segmentation and and transformation for each part is imperfect.
-We often end up with detached parts after the deformation.
-To deal with this, we refine the estimated $\pi_{r}$ and $\pi_{l}$ with the initial segmentation labels in the first phase of optimization using both the Chamfer distance between the transformed $\bar{G}^{t}$ and $G^{t}$, and the rendering loss computed in both the source and target states.
-We define the joint loss as:
+To obtain an initial segmentation and transformations, we leverage TEASER++ [59], a robust pose estimation algorithm.
+For clarity, we first describe the common two-part (root/leaf) scenario (see Sec.
+3.4 for multi-part).
+We apply TEASER++ to the correspondences from Sec 3.2.
+It finds the largest rigid component (inliers) as the root $r$, treating the remaining outliers as the leaf $l$.
+This yields initial transformations $\pi_{r},\pi_{l}$ and part labels $p^{s}$.
+Both the initial geometry from FreeSplatter and the part-pose estimates from TEASER++ are coarse approximations.
+Therefore, a joint optimization is essential to mutually refine the geometry, segmentation, and articulation parameters.
+To do this without the optimization getting stuck in a bad local minimum, we introduce a three-phase staged refinement strategy.
+This coarse-to-fine approach first stabilizes the large-scale rigid motions (Phase 1), then refines the underlying geometry and segmentation (Phase 2), and finally polishes the visual appearance (Phase 3).
+We define our full learnable model and its parameters.
+We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$, where each $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ is augmented with its initial part label $p^{s}$.
+We also treat the rigid transformations $\pi_{r}$ and $\pi_{l}$ as learnable parameters.
+To compensate for non-rigidity and initial estimation errors, we introduce a learnable, per-Gaussian displacement term $\delta\bm{\mu}$.
+To simplify the notation, we define an auxiliary function $f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi)$.
+This function first selects the appropriate rigid transformation $\pi_{p^{s}}$ from the set $\pi$ (using the part label $p^{s}$ contained within $\bar{\mathbf{g}}^{s}_{i}$), applies it, and then adds the displacement $\delta\bm{\mu}_{i}$ to the transformed mean.
+This yields the final deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}_{i}|\bar{\mathbf{g}}^{t}_{i}=f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi),\bar{\mathbf{g}}^{s}_{i}\in\bar{G}^{s}\}$.
+Phase 1: Rigid Pose Refinement The goal of this phase is to stabilize the coarse poses from TEASER++.
+We apply two constraints: the source Gaussian geometry $\bar{G}^{s}$ is frozen, and the non-rigid displacement $\delta\bm{\mu}$ is set to zero.
+Using the fixed initial segmentation $p^{s}$, we only optimize the rigid transformations $\pi=\{\pi_{r},\pi_{l}\}$ by minimizing a joint loss:
 $\mathcal{L}_{\text{joint}}=$ $\mathcal{L}_{\text{CD}}(\bar{G}^{t},G^{t})+$ (4)
 $\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),\mathcal{R}(G^{t}))+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),\mathcal{R}(G^{s}))$
-and our optimization objective is to minimize the joint loss:
+The optimization objective for this phase is:
 $(\pi_{r}^{*},\pi_{l}^{*})=\arg\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (5)
-Note that in this phase, $\bm{\delta}\mu$ is set to zero.
-In the second phase of the optimization, we optimize the Gaussian splats to improve the geometric accuracy, especially to handle occlusion between source and target states.
-We also allow the displacement term $\delta\bm{\mu}$ to be optimized to compensate for non-rigidity caused by imperfect transformation estimation or geometry mismatch, but still penalize the displacement of the Gaussian splats with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$.
-For better optimization stability suggested in ArticulateYourNeRF [7], we update the segmentation labels heuristically.
-Newly split or duplicated points inherit the segmentation labels from their parent points.
-To correct points with incorrect initial segmentation, we check those points with large $\delta\bm{\mu}$, which indicates high non-rigidity, and update their segmentation labels following this rule:
+Phase 2: Geometry and Occlusion Refinement With the rigid poses stabilized in Phase 1, we now address the inaccuracies in the underlying geometry and the initial part segmentation.
+In this phase, we unfreeze the parameters in $\bar{G}^{s}$ and activate the learnable displacement term $\delta\bm{\mu}$.
+This per-Gaussian term aims at compensating for any non-rigid errors, which typically caused by incorrect part segmentations.
+We encourage the displacement to be sparse by penalizing it with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$.
+Crucially, this displacement term also serves as a heuristic to correct mis-segmented points.
+As suggested in [10], we periodically update segmentation labels $p^{s}$.
+We assume points with a large $\delta\bm{\mu}$ are mis-assigned, and we re-assign their label $p$ based on the smallest rigid alignment error:
 $p=\arg\min_{p\in\{r,l\}}\|\bar{\bm{\mu}}^{p}-\bar{\bm{\mu}}^{t}\|_{2}$ (6)
 with $\bar{\bm{\mu}}^{p}$ for the means of $\bar{\mathbf{g}}^{p}=\pi_{p}\circ\bar{\mathbf{g}}^{s},p\in\{r,l\}$.
-And in this phase, our optimization objective is to minimize the joint loss with the regularization term:
+The full optimization objective minimizes the joint loss plus the rigidity penalty, updating all parameters (geometry, transformations, and displacements) simultaneously:
 $\min_{\{\delta\bm{\mu}\},\bar{G}^{s},\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}+\mathcal{L}_{\text{rigid}}$ (7)
-and the update of the segmentation labels is:
-$\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (8)
-In the final phase, we focus on improving visual quality and articulation estimation.
-We update the $\pi_{r}$, $\pi_{l}$, and $\mathbf{c}$ of the Gaussian splats without the displacement term $\delta\bm{\mu}$ for pure rigid transformation supervised by photo loss with the original input images as:
+Phase 3: Pose and Appearance Refinement The previous phases refined the geometry and produced an as-rigid-as-possible part segmentation.
+In this final phase, we lock this geometry (freezing $\bar{G}^{s}$ and setting $\delta\bm{\mu}=\mathbf{0}$) to focus exclusively on finetuning the per-part pose estimation ($\pi_{r},\pi_{l}$) and visual quality (colors $\mathbf{c}$).
+This is supervised purely by the photometric loss against the original input images:
 $\min_{\{\pi_{r},\pi_{l},\mathbf{c}\}}\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{s}),I^{s})+\mathcal{L}_{\text{photo}}(\mathcal{R}(\bar{G}^{t}),I^{t}).$ (8)
 In practice, we use 500 iterations for the first phase, 2500 iterations for the second phase, and 2000 iterations for the final phase.
-After 5000 iterations at total, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
-3.4 Extension to Multiple Moving Parts
+Finally, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
+3.4 Modeling Multiple Moving Parts
 Figure 3: Illustration of the extension to multiple parts. We progressively estimate the segmentation and articulation one part at a time with extra images for new target articulation states.
-For simplicity, we have described our method for two parts, but it can be easily extended to objects with multiple parts as shown in Fig.
+For simplicity, we have described our method for a two-part object.
+Our framework extends to objects with multiple parts by applying it iteratively, as shown in Fig.
 3.
-This extension requires additional input images for the new moving part which are used to initialize the 3D Gaussian splats and camera parameters for each set of images in §3.1.Here we use the Gaussian splats from the second row of input images as source Gaussian and the first row as target Gaussian shown in Fig.
-3.
-We first go through the pipeline with Gaussian deformation in §3.2 and ¢3.3 to model the motion of the first part, which is the light blue arm shown in Fig.
-3.
-After the optimization, we obtain the articulation parameters and segmentation of the first arm.
-Based on this, we use the same source Gaussian together with the Gaussian in the bottom row as target Gaussian to model the second part by repeating deformation §3.2 and joint optimization §3.3.
-Note that there is always only 1 articulated part between source and target Gaussian.
+This requires a common source image set ($\mathcal{I}^{s}$) and additional target sets ($\mathcal{I}^{t1},\mathcal{I}^{t2},\dots$), one for each moving part.
+We first run the full pipeline on the $(\mathcal{I}^{s},\mathcal{I}^{t1})$ pair to model the first part.
+We then repeat the process on the $(\mathcal{I}^{s},\mathcal{I}^{t2})$ pair to model the second part, and so on.
+The key assumption is that each source-target pair isolates the motion of only one part.
 4 Experiments
-[…]
-Motion Ang Err $\downarrow$ AGS[10] GT 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
-AYN[7] GT 52.986 60.834 2.105 42.101 81.157 88.316 74.938 24.850 failed 66.718
-AGS[10] VGGT 49.001 66.191 82.359 51.069 21.378 40.371 288.473 83.680 67.020 71.801
-Ours 0.095 0.914 1.020 2.365 0.839 1.337 2.905 1.737 0.878 0.525
-Pos Err $\downarrow$ AGS[10] GT 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
-AYN[7] GT 0.021 0.073 0.033 2.116 0.221 0.231 0.916 0.334 - -
-AGS[10] VGGT 0.455 0.254 0.033 0.775 0.741 0.090 0.412 0.513 - -
-Ours 0.180 0.007 0.106 0.042 0.058 0.196 1.042 0.157 - -
-Motion Dist $\downarrow$ AGS[10] GT 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
-AYN[7] GT 122.730 562.642 10.115 20.583 111.252 118.564 50.044 81.205 failed 0.373
-AGS[10] VGGT 104.417 54.799 43.799 45.248 27.312 57.388 53.142 34.177 0.202 0.297
-Ours 0.861 2.279 0.788 2.654 1.516 0.930 2.227 2.926 0.017 0.118
-Rendering PSNR $\uparrow$ AGS[10] GT 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
-AYN[7] GT 10.827 14.631 13.411 15.540 14.675 15.826 16.801 15.784 failed 14.018
-AGS[10] VGGT 10.740 16.403 17.737 15.123 18.755 14.866 12.990 16.082 18.268 18.483
-Ours 24.659 26.920 26.491 23.434 28.823 27.900 27.016 25.408 29.041 25.493
-SSIM $\uparrow$ AGS[10] GT 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
-AYN[7] GT 0.718 0.893 0.852 0.816 0.904 0.881 0.855 0.889 failed 0.814
-AGS[10] VGGT 0.693 0.861 0.883 0.801 0.910 0.890 0.821 0.878 0.905 0.874
-Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
-Geometry CD $\downarrow$ AGS[10] GT 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
-AYN[7] GT 677.761 164.321 124.728 335.108 258.443 125.208 740.898 1575.325 failed 62.890
-AGS[10] VGGT 476.935 48.592 77.646 26.251 21.667 112.274 632.828 24.736 26.769 15.599
-Ours 0.856 2.051 0.610 6.449 0.264 1.216 8.935 1.615 1.788 6.443
-Table 1: Comparisons to the baseline methods with 4-view images from each articulation states (source and target). Objects with $\dagger$ are with prismatic joints.
+Corr. Oven Storage USB Laptop
+GaussReg Refer to caption Refer to caption Refer to caption Refer to caption
+FM-Dense Refer to caption Refer to caption Refer to caption Refer to caption
+Ours Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 4: Qualitative analysis of correspondence methods.
+We visualize the part segmentations computed by our TEASER [59] solver, which is fed correspondences from each method.
+Inliers for the two main segmented parts are colored blue and orange; outliers and non-correspondences are shown in gray.
+Better view in color and zoom in.
 4.1 Experimental Setup
@@
 For PSNR, SSIM and CD, we report the average values of the two articulation states.
-Baselines.
-To the best of our knowledge, no existing work addresses self-supervised articulated object modeling from few-shot unposed images.
-Nevertheless, we can evaluate our approach by comparing it with state-of-the-art methods that utilize posed images for articulated object modeling, specifically ArticulatedGS (AGS) [10].
-For fair comparison, we employ VGGT [44] to estimate camera poses for the few-shot input images and point clouds initialization.
-We thus compare our method against two baselines: AGS with ground truth poses and AGS with VGGT estimated poses plus VGGT estimated point clouds, using 4 images per articulation state.
-[…]
-[…]
+GT AGS AGS-Full Ours
+FM-Dense Ours
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption failed Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption failed Refer to caption
 Figure 5: Qualitative evaluation for novel view synthesis in target state.
@@
 In the meanwhile, our method demonstrates similar rendering quality compared to AGS-Full, which is trained with 100 images per articulation state with ground truth camera poses.
+Baselines.
+To our knowledge, no existing work addresses self-supervised articulated object modeling from sparse, unposed images.
+Our evaluation is therefore twofold: a component-wise analysis to validate our correspondence method and a full-pipeline comparison against state-of-the-art methods.
+For the component analysis, we fix our pipeline and swap our correspondence module with two alternatives: (1) GaussReg [3], a pre-trained rigid GS registration method that provides 1024 sparse keypoint correspondences, and (2) FM-Dense [2], a pre-trained non-rigid matching network, to which we feed 20480 downsampled points from our GS to obtain dense correspondences.
+For the full-pipeline comparison, we evaluate against ArticulatedGS (AGS) [16] with the same number of images ($K=4$) and ground truth, aligned camera parameters.
+Besides, we use the VGGT [53] reconstructed point clouds to initialize the AGS for better reconstruction.
+We will make our code and models publicly available.
 4.2 Comparison with Baselines
-Table 1 presents the quantitative comparison between our method and the baselines.
-The results clearly show that existing methods cannot handle the few-shot unposed setting—all three methods perform catastrophically across all metrics.
-Despite using VGGT-estimated point clouds for initialization, AGS-VGGT still fails to produce meaningful reconstructions.
-Our method, however, consistently outperforms the baselines across all metrics and object types.
-The baseline methods exhibit very high Chamfer Distance errors, indicating poor geometry reconstruction quality even at the source state.
-The poor performance stems from several fundamental limitations: limited input views severely hinder baseline reconstruction capabilities.
-Since these baseline methods heavily rely on precise 3D reconstruction and camera poses to identify articulated parts, the poor geometry reconstruction quality inevitably leads to inaccurate articulation parameter estimation and degraded novel articulation synthesis quality.
-As for the AYN method, it raises an error when running on the "Blade" object as it failed to initialize part segmentation due to poor reconstruction quality.
-Despite outperforming the baselines, certain objects remain challenging for our method.
-For instance, the part segmentation of the "Blade" object shows limited accuracy, as illustrated in Fig.
-5.
-We observe that only the blade tip and the blue handle are properly segmented, while the main blade body is classified as the root part.
-This occurs because the blade’s movement is too small, the pipeline could produce similar visual results during rendering by just moving the blade tip and the blue handle.
-Additionally, our method experiences performance degradation when dealing with thin articulated parts, like the lid of "Washer" and the door of "Oven".
-Since reconstructing thin objects is inherently challenging, the resulting geometry errors create larger margins for articulation estimation errors.
-Besides, objects with large occluded parts are difficult to to maintain consistency in the reconstructed geometry between source and target states.
-One example would be the USB, which has a large occluded part at target state.
-Such geometry inconsistency results in performance degradation.
-We further examine the baseline failures through visual analysis.
-Figure 4 shows view synthesis results, where baseline methods generate heavily corrupted renderings with substantial noise—the objects become barely recognizable.
-Surprisingly, our method matches the quality of AGS-Full, which uses 100 images per articulation state with ground truth poses.
-Figure 5 displays the 3D Gaussian point distributions with part segmentation.
-The baselines produce scattered floating points and fragmented geometries, while our method creates coherent 3D structures with proper part segmentation and articulation parameter estimation.
-[…]
-rowsep=0pt,colsep=1pt RGB input AGS-GT AGS-VGGT AGS-Full Ours GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Visualization of 3D Gaussian points with part-level segmentation and articulation axes. Ground truth axises are shown in green, predictions in red. Better view in color and zoom in.
-[…]
-AGS GT 34.914 0.988 1.279 0.101
-VGGT 28.860 0.975 1.367 1.623
-AYN GT 32.219 0.962 0.946 0.654
-Ours - 28.870 0.975 1.399 1.760
-Table 2: Comparative analysis for 16 input views per articulation state.
-Figure 6: PSNR w.r.t. number of views per articulation.
-[…]
-rowsep=1pt,colsep=1pt 4-view 8-view 16-view AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AYN-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 7: Qualitative evaluation for the effect of the number of input images.
-We show novel view synthesis for source state (left) and target state (right) at each settings.
-From top to bottom: AGS-GT, AYN-GT, AGS-VGGT, Ours.
-[…]
-rowsep=0pt,colsep=1pt $k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$ Refer to caption Refer to caption Refer to caption Coarse segmentation at source state Refer to caption Refer to caption Refer to caption Deformation Overlay of Gaussian Points
+Quantitative Evaluation
+Quantitative results are presented in Tab. 1.
+We first compare our method against ArticulatedGS (AGS) [16].
+The results show that AGS, even with GT aligned camera poses and VGGT initialization, fails to converge in a sparse 4-view setting.
+This is because its optimization, designed for dense data, introduces significant noise from sparse inputs, and the limited views are insufficient for its motion estimation to converge.
+These factors lead to catastrophic errors, demonstrating that existing SOTA methods are not robust to this sparse-input scenario.
+Our qualitative analysis in Fig.
+4 visually confirms our quantitative findings by showing the output of the TEASER [59] solver for each baseline (inliers in blue/orange, outliers in gray).
+The GaussReg [3] output is entirely gray, confirming its 1024 sparse, noisy correspondences are insufficient to initialize segmentation.
+FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part.
+In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers.
+A small volume of outliers is visible at part boundaries (e.g., USB, Storage), which are refined in the joint optimization step.
+RGB input AGS-GT AGS-Full FM-Dense Ours GT
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 6: Visualization of 3D Gaussian points with part-level segmentation and articulation axes.
+Ground truth axises are shown in green, predictions in red.
+AGS-Full denotes the AGS pipeline with dense and posed images input.
+Better view in color and zoom in.
+$k_{0}=10^{-5}$ $k_{0}=10^{-8}$ $k_{0}=10^{-11}$
+Refer to caption Refer to caption Refer to caption
+Coarse segmentation at source state
+Refer to caption Refer to caption Refer to caption
+Deformation Overlay of Gaussian Points
 Figure 7: Ablation study for the choice of $k_{0}$ for the deformation field.
@@
 For deformation overlay, red points denote the deformed Gaussian points $\hat{G}^{t}$, and green points denote the Gaussian points at target state $\hat{G}^{t}$.
-[…]
-rowsep=0pt,colsep=1pt RGB input Novel articulation synthesis Part-level segmentation Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+RGB input Novel articulation synthesis Part-level seg.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 8: Real-world objects.
@@
 The green lines denote the estimated articulation axes.
+Qualitative Evaluation
+We provide a qualitative analysis of the correspondence methods in Fig.
+4.
+This figure visualizes the output of the TEASER [59] solver when fed correspondences from each baseline.
+Inliers for the two main parts are colored blue and orange, while outliers and non-correspondences are gray.
+The results visually confirm our quantitative findings.
+The GaussReg [3] baseline fails completely; its point cloud is rendered entirely gray, confirming that its 1024 sparse and noisy keypoint correspondences are insufficient to initialize segmentation.
+FM-Dense [2] proves highly unstable; the visualization shows most of its 20k correspondences are rejected as outliers, and even when a solution is found (e.g., Laptop), segmentation is incorrect, with inliers (blue/orange) assigned to the same part.
+In contrast, our method produces a dense, reliable map with a high percentage of correctly segmented inliers onto the distinct rigid parts.
+A small volume of outliers (gray points) is visible at the part boundaries on objects like the USB and Storage, which will can be further optimized by the joint optimization step.
+Further qualitative evidence is provided in Fig.
+6, which visualizes the final part segmentation and estimated joint axes.
+For this comparison, we include AGS [16] (using its full, dense-view and posed-image configuration) as a state-of-the-art upper bound.
+As the figure illustrates, our method’s joint parameter estimation is visually on par with the SOTA AGS baseline, despite our significantly more challenging sparse and unposed input.
+In contrast, the FM-Dense [2] baseline demonstrates degraded performance, resulting in clear errors in both segmentation and joint estimation.
+This visualization further highlights the robustness and accuracy of our full pipeline.
+We also provide the qualitative evaluation for novel articulation synthesis in Fig.
+5.
+For more visualizations, please refer to the supplementary.
 4.3 Ablation Studies
-Effect of the number of input images.
-With just 4-view images, baseline methods struggle to reconstruct objects even when in static state.
-We investigate how the number of input views affects performance by varying the views per articulation state.
-Fig.
-6 shows that PSNR for baseline methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses.
-Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation shows in Fig.
-7.
-Tab.
-2 shows the comparison at 16 images per state, where AGS with VGGT poses gets close to our performance.
-However, AGS still encounter reconstruction failures even with 32 images per state when VGGT provides less accurate poses.
-We suspect that AGS is more sensitive to pose estimation errors as it heavily depends on 2D image information.
-Our approach leverages FreeSplatter’s 3D priors and 3D-guided deformation, making it robust to pose estimation variations when poses stay consistent relative to the reconstructed splats.
 Effect of $k_{0}$ for deformation field
@@
 To analyze its effect, we conduct experiments with different $k_{0}$ values, with qualitative results shown in Fig.
-8.
+7.
 A smaller $k_{0}$ enforces near-rigid transformations across the whole object, leading to under-deformation and failure to reach the target state (e.g., right side of the second row).
@@
 Empirically, we find that setting $k_{0}$ between $10^{-7}$ and $10^{-9}$ achieves a good balance, preserving rigidity while allowing sufficient flexibility for deformation.
-4.4 Real-world Experiments
-To demonstrate the generalization capability of our approach, we conduct experiments on real-world articulated objects.
-We evaluate our method on four different objects: a bottle, a projector, a drawer, and a lego toy, collecting four input images for each articulation state.
-Prior to processing, we employ the Segment Anything Model (SAM) [36] to remove background from the input images.
-For objects with multiple articulated parts, such as the lego toy with two movable arms, we collect three sets of input images per articulation state as explained in Sec.
+Real-world Experiments
+To demonstrate the generalization of our approach, we conduct experiments on four real-world objects (a bottle, projector, drawer, and lego toy), capturing four input images per articulation state.
+We employ the Segment Anything Model (SAM) [43] to remove backgrounds prior to processing.
+For multi-part objects, like the lego toy, we capture three image sets (one per state) as detailed in Sec.
 3.4.
-The experimental results, presented in Fig.
-9, demonstrate that our method successfully generates seamless reconstruction results for real-world articulated objects, validating its practical applicability beyond synthetic datasets.
+The results in Fig.
+8 show that our method successfully generates seamless reconstructions for these real-world objects, validating its practical applicability.
 5 Conclusion
-We introduced PAOLI, a self-supervised framework for learning articulated object representations from sparse, unposed images.
-Unlike prior methods that require dense multi-view data and known camera poses, PAOLI operates with only a few views per articulation and no camera supervision.
-By combining sparse-view 3D reconstruction, a deformation field for cross-pose correspondences, and a progressive disentanglement strategy, our method jointly optimizes geometry, appearance, and kinematics while robustly separating camera and object motion.
-Our work assumes articulated objects with rigid parts, cannot handle non-rigid deformations.
-Handling such transformations requires more complex deformation models.
-Second, we assume that the observations from two different poses of the same object are available.
-It would be valuable to pursue an active vision approach that allows the system to interact with the object while learning its articulated structure, which could improve robustness and generalization (e.g., [34]).
-Additionally, our current method relies on comprehensive object coverage with four input views per articulation state; insufficient coverage may degrade reconstruction quality and lead to failure.
-Integrating diffusion models to synthesize additional viewpoints could offer a promising solution for scenarios with limited view coverage.
-Finally, temporal information could be leveraged to improve the understanding of articulated motion dynamics.
+We have introduced PAOLI, the first framework to successfully learn high-fidelity articulated object representations from sparse, unposed images.
+We identified that the central challenge in this setting is not just reconstruction, but the 3D-to-3D correspondence problem between unaligned 3D models.
+Our solution is a test-time, optimization-based deformation field that finds dense and robust correspondences by leveraging both 3D geometry and 2D rendering losses.
+Our experiments show this approach is highly effective, whereas prominent baselines—including sparse keypoint methods and pre-trained dense networks—fail to solve the problem.
+Our full pipeline successfully recovers high-fidelity geometry, appearance, and kinematics, demonstrating a robust path forward for articulated object learning in practical, unposed scenarios.
+While effective, our method assumes rigid parts and requires at least two distinct articulation states with adequate view coverage.
+Furthermore, our alignment may fail if the initial, arbitrary coordinate systems of the 3D GS reconstructions are extremely misaligned.
+More detailed discussion on limitations can be found in the supplementary.
+[…]
 Supplementary Material
 6 Supplementary Material
-6.1 Detailed Discussion of Related Work
-Here we provide more detailed discussion of the most related articulated object learning work including, PARIS [22], AYN [7] and AGS [10].
-As explained in the submission, these techniques assume dense views of the object across two articulation state along with the camera information, unlike our method which is pose-free and works with sparse views.
-Below, we explain the differences in methodology in more detail.
-All the methods including ours uses the synthesis by analysis approach that deforms the object shape in the source articulation state to the target articulation state while learning the geometry, appearance and part segmentation as the intermediate representations.
-PARIS jointly optimizes the geometry, appearance and articulation parameters from the input images of both source and target states with photometric loss.
-It employs two separate NeRFs to model ‘static’ and ‘dynamic’ parts of the object.
-Although the joint optimization is straightforward, it is unstable and fails to obtain accurate results when the initialization of these parameters are noisy, as reported in both AYN and AGS.
-In addition, PARIS is limited to modeling one moving part only.
-Instead, both AYN and AGS follows a multi-step approach, hence reconstructs the geometry and appearance of the object first in the source state.
-Once the initial geometry and appearance are learned, they use a decoupled optimization for part segmentation and motion parameter estimation that provides more stable optimization.
-AYN first extracts evenly spaced point cloud from NeRF scans, and then projects the point clouds back on image space.
-By comparing the difference in the alpha channel between the projected images and the target state images, AYN initializes the part segmentation label.
-With the initialized labels, it optimizes the part motion parameter so that the projected images with transformed part-points match the image of corresponding target images in alpha channel measured by 2D chamfer distance.
-Later, once the motion parameters are fixed, and AYN learns a segmentation field to refine the segmentation by minimizing the photo metric loss between rendered images and target images.
-Similarly, AGS tackles the instability problem with multi-step approach.
-To first obtain the segmentation labels for gaussians, by employing a deformation field to learn the displacement for each Gaussian and rotation to correct the quaternions for color rendering.
-That’s $\hat{\bm{\mu}}=\bm{\mu}+\delta\bm{\mu}$, $\hat{\bm{q}}=\bm{q}\otimes\delta\bm{q}$.
-This steps aims to obtain the deformed 3D Gaussian that could render similar images in the target states with the corresponding camera poses.
-The deformation is supervised by photometric loss and a rigidity regularization term.
-Since the camera poses are aligned with the root part of the objects, the gaussian points in the root part will be static with small displacement $\delta\bm{\mu}$.
-AGS assigns Gaussians with small displacement to the static part, and the ones with large deformations to the moving part.
-After freezing their segmentation labels, it estimates a rigid transformation for the dynamic part with chamfer distance w.r.t.
-the deformed Gaussian (by deformation field) and photometric loss w.r.t.
-the images in the target states.
-However, both AYN and AGS rely on known camera poses that are aligned at the root part between source and target states.
-This requirement prevents them from initializing part segmentation when camera poses are unknown across different articulation states.
-Unlike previous methods that identify parts based on rendering differences (which only relies on 2D information from images) and assume a static root part, our method establishes correspondence across two articulation states directly in 3D space by deforming the 3D Gaussians from the source to the target state.
-Since $G^{s}$ and $\hat{G}^{t}$ represent the same set of Gaussians with $\hat{G}^{t}$ being the transformed version of $G^{s}$, we can naturally establish matching pairs between them.
-The deformation is supervised by 3D chamfer distance and photometric loss, ensuring that $\hat{G}^{t}$ closely approximates $G^{t}$.
-This makes the matching pairs from $G^{s}$ to $\hat{G}^{t}$ effectively equivalent to matching pairs between $G^{s}$ and $G^{t}$.
-With these correspondences established, we can separate parts and estimate per-part transformations using TEASER for initialization.
-To further improve rigidity during deformation and enhance performance, we employ positional encoding on the Gaussian means before feeding them to the deformation network.
-With the initialization, we can now continue for fine-level optimization.
-For stability purpose, we also decouple the updates for part segmentation from the joint optimization.
-We split the optimization into 3 phases.
-The first phase we try to refine the coarse motion estimation from TEASER.
-In the second phase, we update part segmentation based on heuristic rules explained in §3.3 in the main paper to refine the segmentation.
-At the last phase, we optimize the appearance to improve the rendering quality.
+6.1 Implementation Details
+Deformation Network.
+The deformation network $\mathcal{F}_{\text{deform}}$, introduced in Sec.
+3.2, is a 3-layer MLP with 128-dimensional hidden layers and ReLU activations.
+The network outputs a rigid transformation where rotation is parameterized as Euler angles.
+Computation Efficiency
+Our full pipeline runs in approximately 15 minutes per object on a system with an AMD Zen3 7950X CPU and an NVIDIA RTX 4090 GPU. The runtime breaks down as follows:
+  • Step 1: $\sim$1 minute for initial reconstruction and camera pose refinement.
+  • Step 2: $\sim$4 minutes for deformation field optimization.
+  • Step 3: $\sim$10 minutes, which includes a few seconds for the TEASER++ solver and the remainder for joint refinement.
+The peak memory usage is approximately 15 GB of VRAM during the Step 3.
+6.2 Ablation Studies
+Full Pipeline w.o. Cam. Opt. w.o. $L_{CD}$ w.o. $L_{photo}$
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 9: Visualization of the results after TEASER initialization in Step 3. The main impact of these ablated components occurs before this initialization step. Better view in color and zoom in.
+Impact of Key Pipeline Components.
+Table 2: Ablation study on the key components of our pipeline.
+We report the average metrics across all objects.
+Removing any component results in performance degradation, highlighting their importance.
+Components Metrics
+cam_opt $\mathcal{L}_{\text{CD}}$ $\mathcal{L}_{\text{photo}}$ Ang Err $\downarrow$ Pos Err $\downarrow$ Motion Dist $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$
+$\checkmark$ $\checkmark$ $\checkmark$ 1.160 0.105 1.318 26.573 0.953 2.272
+[…]
+[…]
+[…]
+We conducted an ablation study to analyze the impact of key components in our pipeline: camera pose refinement in step 1 and the losses used for deformation network optimization ($\mathcal{L}_{\text{CD}}$ and $\mathcal{L}_{\text{photo}}$) in step 2.
+As shown in Tab.
+2, removing any of these components leads to a degradation in performance.
+These components are crucial for initializing the part segmentation and pose estimation in Step 3.
+To understand the performance drop, we examined the initialization results.
+As illustrated in Fig.
+9, omitting these components leads to poorer initializations, which subsequently hampers the final joint optimization stage.
+This demonstrates that each component plays a vital role in the robustness of our framework.
+Effectiveness of the Joint Optimization
+The effectiveness of our joint optimization process is demonstrated in Fig.
+10.
+The figure shows how the process corrects initial errors: despite dense matching, the state after TEASER initialization still has spatial discrepancies between parts.
+Phase 1 recovers connectivity, Phase 2 refines part segmentation, and Phase 3 produces the final, rigid, joint-optimized output.
+TEASER init. Phase 1 Phase 2 Phase 3
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 10: Visualization of the effectiveness of the joint optimization process, showing the progression from the initial state to the final refined output.
+Effect of the number of input images.
+With just 4-view images, traditional articulated object modelling methods struggle to reconstruct objects even when in static state.
+We investigate how the number of input views affects performance by varying the views per articulation state.
+Fig.
+11 shows the rendering quality for ArticulatedGS [16] methods improves with more input images, saturating after 16 images (VGGT runs out of memory with 64 images on our 48GB VRAM GPU), even with VGGT-estimated poses.
+Our method matches the saturated performance without ground truth poses with only 4 input images and stays consistent across different view counts with qualitative evaluation.
+More Qualitative Results
+We provide additional qualitative results for correspondences and segmentation in Fig.
+13 and Fig.
+15, respectively.
+Besides, correspondences visualization on the real world images are also shown in Fig.
+14.
+6.3 Failure Cases
+Extreme Initial Misalignment.
+We demonstrate some failure cases in Fig.
+12.
+Our method relies on FreeSplatter [62] for initial 3D reconstruction, which establishes a coordinate system based on the first input image of a set.
+If the first images from the source and target sets are captured from drastically different viewpoints (e.g., front vs.
+back), the resulting 3D models can be severely misaligned.
+This extreme misalignment can cause our deformation network to converge to a poor local minimum, especially for objects with geometric symmetries.
+For instance, as shown with the real-world storage drawer in Fig.
+7, providing opposing side views as initial frames causes the network to favor a non-rigid deformation of the drawer rather than finding the correct global alignment.
+To mitigate this, we recommend selecting initial views from similar perspectives for each image set.
+4-view 8-view 16-view
+AGS-GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+AGS-VGGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 11: Qualitative evaluation for the effect of the number of input images.
+We show novel view synthesis for source state (left) and target state (right) at each settings.
+From top to bottom: AGS-GT, AGS-VGGT, Ours.
+Input reference view Render deformation target states 3D overlay
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 12: Visualization of the failure cases in extreme misalignment case (top) and thin structures with large motion (bottom).
+The 3D overlay shows the itermediate state during the optimization of deformation field, the red is the deformed source state, the green is the target state.
+Thin Structures with Large Motion.
+The framework can struggle with objects that have very thin parts undergoing large articulation.
+A thin structure is represented by a smaller volume of Gaussian splats, which provides weaker rigidity constraints during the deformation step.
+When combined with a large motion, the deformation field may fail to preserve the part’s rigidity, leading to inaccurate correspondence and subsequent errors in joint estimation.
+An example of this can be seen in Fig.
+12.
 6.4 Evaluation Metrics
@@
 For prismatic joints, we compute the distance between the predicted translation magnitude and the ground truth translation magnitude.
-6.3 Additional Visualizations
-Source Input Target Input Novel View/Articulation Synthesis
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Ours FM-Dense [2] GaussReg [3] Ours FM-Dense [2] GaussReg [3]
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 13: Visualization of correspondences for our methods and baselines.
+Inliers for the two main rigid parts (from TEASER solver) are colored green and orange.
+For GaussReg [3], outliers are shown in black.
+Since our method and FM-Dense [2] produce dense correspondences, we visualize a sample of up to 2000 inliers per part.
+Better view in color and zoom in.
+Source Input Target Input Correspondences
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 14: Visualization of correspondences for our methods and baselines on real-world data. Better view in color and zoom in.
+Source/Tgt Novel View/Articulation Synthesis Seg.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 15: Additional qualitative results for novel view and articulation synthesis.
+For each object, we show one source and one target input view, followed by a sequence of renderings from our model with novel interpolated unseen articulation states.
+We also show the segmentation with estiamted asxis at the end of each row.